OpenStudy (anonymous):

i need to "solve for the indicatd variable if the line throught the two given points has the given slope." (2,b) and (-1, 4b), m=-2 How do I begin this problem? the b variable is throwing me off.

OpenStudy (anonymous):

what formula do i use?

OpenStudy (anonymous):

Use m = (y2 - y1) / (x2 - x1)

OpenStudy (anonymous):

thank you...

OpenStudy (blacksteel):

LA has it right. The formula for slope is change in rise (y) over change in run (x). We know that the points (2,b) and (-1,4b) are on our line, and we know that between these two points, the change in rise is b-4b = -3b and the change in run is 2 - (-1) = 3.Thus, our slope m = -3b/3. Then m = -b = -2, so b = 2.